# Structured Learning of Compositional Sequential Interventions

### v1

@@ 6 Conclusion
We introduced an approach for predicting potential outcomes with a careful accounting of when extrapolation to unseen sequences are warranted. Lessons. Embeddings are important given sparse categorical sequential data see e.g. (Vaswani et al., 2017). However, large combinatorial interventional spaces benefit from models that more carefully lay down conditions for identification. Limitations. Unlike the traditional synthetic control literature Abadie (2021), we assume a model for time effects based on autoregression and truncated or parametric temporal-intervention interactions. While it is possible to empirically evaluate the prediction abilities of the model using a validation sample, high-stakes applications (such as major interventions to counteract the effect of a pandemic) should take into consideration that uncontrolled distribution shifts may take place, and careful modeling of such shifts should be added to any analytical pipeline to avoid damaging societal implications. Future Work. Besides allowing for an explicit parameterization of drifts, an explicit account for modeling when unmeasured confounding may have been presented in the selection of past interventions is necessary to increase the applicability of the method to more sources of data.
@@ Acknowledgements
We thank Zhenwen Dai and Georges Dupret for many helpful comments and suggestions. JY and RS were partially funded by the EPSRC Open Fellowsship EP/W024330/1. RS acknowledges further funding from the EPSRC AI Hub for Causality in Healthcare AI with Real Data, EP/Y028856/1.
@@ Appendix A General Background
Much of traditional causal inference aimed at estimating an outcome of a fixed treatment that does not vary over time Imbens and Rubin [2015]. However, many causal questions involves treatments that vary over time. For example, we might be interested in estimating the causal effects of medical treatments, lifestyle habits, employment status, marital status, occupation exposures, etc. Because these treatments may take different values for a single individual over time, hence they are often refereed to as time-varying treatments [Hernán and J, 2020].
In many designs, it is assumed that a time-fixed treatment applied to individuals in our population happens at the same time (i.e. there is a shared global clock for everyone). A more relaxed setting (and more practical, as described in this paper) is that (categorical) interventions can happen to different individuals at different time steps, potentially resulting in no two identically distributed time-series. The purpose now is to estimate the causal effect or predict expected outcomes under time-varying treatments as a time-series, rather than estimating missing potential outcome snapshots at fixed time interval that has passed already.
One typical way of designing experiments under a binary treatment (00 or $1$) [Robins, 1986, 1997] of time-varying treatments is to have $\bar{d}_{0}=(0,0,\cdots,0)$ (never treated) and $\bar{d}_{1}=(1,1,\cdots,1)$ (always treated) as the two options in the design space. This allows us to answer what the average expected outcome contrasted between these two options. A fine-grained estimation at every time step will require further levels of contrast under two regimes, such as $d=(0,1,\cdots,1)$ and $d=(1,0,\cdots,0)$. However, the number of combinations grows exponentially on $T$, the number of time steps, even when only binary treatments are at the stake. In this paper, we focus on static treatment strategies where the treatment is randomly assigned (also referred to as a sequential randomized experiment), although the next intervention in a sequential randomised experiment can depend on previous step intervention, as long as there is no influence from unobserved confounders.
Conditional ignorability holds when we are able to block unmeasured confounders using observable confounders only. Similarly, in a time-varying setting, we achieve sequential conditional ignorability by conditioning on the time-varying covariates, including past actions (Hernán and J [2020], and Chapter 4 of Pearl [2009]). In our manuscript, we will have little to say about the uses of conditional sequential ignorability and other adjustment techniques for observational studies, as these ideas are well-understood and can be applied on top of our main results. Hence, we decided to focus on treatments that act as-if randomized so that we can explore in more detail our contributions.
@@ Appendix B Proofs
Let $f(n,d^{1:T},x^{1:T},z):=f_{d}^{\mathsf{T}}(d^{1:T})f_{nxz}(n,x^{1:T},z)$, where function sequences $f_{d}(d^{1:T})$ and $f_{nxz}(n,x^{1:T},z)$ are defined for all $T\in\mathbb{N}^{+}$ and have codomain $\mathbb{R}^{r_{g}}$. Furthermore, define $f_{dl}(d^{1:T})$ as in Eq. (4) with $r_{h}=1$, and the $l$-th entry of $f_{nxz}(n,x^{1:T},z)$ as $f_{nxzl}(n,x^{1:T},z):=u_{l}(n)^{\mathsf{T}}v_{l}(x^{1:T},z)$, each of the latter having codomain $\mathbb{R}^{r_{s^{\prime}}}$ and where $v_{l}(\cdot,\cdot)$ is defined recursively for any $T$, similarly $g_{l}(\cdot,\cdot,\cdot)$. Then there exists some integer $r$ and three functions $a,b$ and $c$ with codomain $\mathbb{R}^{r}$ so that $f(n,d^{1:T},x^{1:T},z)=\sum_{l=1}^{r}a_{l}(n)\times b_{l}(x^{1:T},z)\times c_{l}(d^{1:T},T)$. Moreover $c_{l}(d^{1:T},T)=\prod_{t=1}^{T}m_{l}(d^{t},t,T)$ for functions $m_{l}:\mathbb{N}^{3}\rightarrow\mathbb{R}$.
@@
$\begin{array}[]{rcl}f_{d}^{\mathsf{T}}(d^{1:T})f_{nxz}(n,x^{1:T},z)&=&\sum_{l^{\prime}=1}^{r_{s}}f_{dl^{\prime}}(d^{1:T})f_{nxzl^{\prime}}(n,x^{1:T},z)\\ \\ &=&\sum_{l^{\prime}=1}^{r_{s}}f_{dl^{\prime}}(d^{1:T})\sum_{l^{\prime\prime}}^{r_{s^{\prime}}}u_{l^{\prime}l^{\prime\prime}}(n)v_{l^{\prime}l^{\prime\prime}}(x^{1:T},z)\\ \\ &=&\sum_{l^{\prime\prime}=1}^{r_{s^{\prime}}}\sum_{l^{\prime}=1}^{r_{s}}u_{l^{\prime}l^{\prime\prime}}(n)v_{l^{\prime}l^{\prime\prime}}(x^{1:T},z)f_{dl^{\prime}}(d^{1:T})\\ \\ &=&\sum_{l=1}^{r}a_{l}(n)b_{l}(x^{1:T},z)c_{l}(d^{1:T}),\end{array}$
@@
Finally, each $c_{l}(d^{1:T})$ is by definition some $f_{dl^{\prime}}(d^{1:T})$. From Eq. (4) and the assumptions, $f_{dl^{\prime}}(d^{1:T})=g_{l^{\prime}}(d^{1:T},T,T)$, and $g_{l^{\prime}}(d^{1:t},t^{\prime},t):=h_{l^{\prime}}(g_{l}(d^{1:t^{\prime}-1},t^{\prime}-1,t),d^{t^{\prime}},t^{\prime},t)$ for $t^{\prime}>0$ and $d>0$, with otherwise $h_{l^{\prime}_{1}}(g_{l^{\prime}}(d^{1:t},t^{\prime}-1,t))h_{l^{\prime}_{2}}(d^{t},t^{\prime},t)$. For $t^{\prime}=0$, $h_{l^{\prime}}(v,d,t^{\prime},t)\equiv 1$, and for $t^{\prime}>0$ and $d=0$, $h_{l^{\prime}}(v,d,t^{\prime},t)\equiv g_{l^{\prime}}(d^{1:t^{\prime}-1},t^{\prime}-1,t)$. By recursive application, this implies $f_{dl^{\prime}}(d^{1:T})=\prod_{t^{\prime}=1:d^{t^{\prime}}>0}^{T}h_{l^{\prime}_{2}}(d^{t^{\prime}},t^{\prime},T)$ and the result follows. $\Box$
@@
Suppose we have a measurable function $f:\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1}\times\bigcup_{t=1}^{T}\mathcal{D}^{t}\times\mathcal{Z}\rightarrow\mathbb{R}^{d}$, $t\leq T$, where $\mathcal{X}^{1:t-1}$ and $\mathcal{Z}$ are compact, $\mathcal{D}^{t}$ is finite. For $d=1,\cdots,d_{max}$, let $m_{d}:\bigcup_{t=1}^{T}\mathcal{D}^{t}\rightarrow\{1,\cdots,t\}$ with $m_{d}(\mathbf{d})=t\text{ if }\exists t\;s.t.\;\mathbf{d}_{t}=d\;\text{else }0$. There exist $L=O(T^{d_{max}+1})$ such that for any $\epsilon$, there are measurable functions $\phi_{l}:\mathcal{X}^{1:t-1}\times\mathcal{Z}\rightarrow\mathbb{R}^{d}$ and $\psi_{ld^{\prime}}:\mathcal{D}^{t}\rightarrow\mathbb{R}$ and real numbers $\beta_{l}$ for $l=1,\cdots,r$ and $d=1,\cdots,d_{max}$, $\hat{f}(\mathbf{x},\mathbf{d},z)=\sum_{l=1}^{L}\phi_{l}(\mathbf{x},z)\beta_{l}\prod_{d=1}^{d_{max}}\psi_{ld^{\prime}}(m_{d^{\prime}}(\mathbf{d}))$ such that
$\text{sup}_{\mathbf{x}\in\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1},\mathbf{d}\in\bigcup_{t=1}^{T}\mathcal{D}^{t},z\in\mathcal{Z}}\left|f(\mathbf{x},\mathbf{d},z)-\hat{f}(\mathbf{x},\mathbf{d},z)\right|\leq\epsilon$ (11)
Proof. First note that we can construct an equivalent definition of $D^{t}$ for some fixed $t$ as follows. We for now call the new object $\bar{D}^{t}$.
$\bar{\mathcal{D}}^{t}=\left\{\mathbf{v}\in\mathbb{R}^{d_{max}}\bigg{|}v_{d}\in\{0,\cdots,t\}\forall d;\;v_{d^{\prime}}\neq v_{d}\text{ whenever $d^{\prime}\neq d$ and $v_{d}\neq 0$}\right\}$ (12)
Note that there is a one-to-one correspondence between $\bar{\mathcal{D}}^{t}$ and $\mathcal{D}^{t}$, and denote the bijection connecting them by $g:\bar{\mathcal{D}}^{t}\rightarrow\mathcal{D}^{t}$. Since both sets are finite, we can approximate $f$ with $\hat{f}$, if and only if we can approximate $f\circ i\times g\times i$ by $\hat{f}\circ i\times g\times i$. So wlog just consider $\bar{\mathcal{D}}^{t}$ as $\mathcal{D}^{t}$.
@@
Now note, fixing the ground truth function $f$ at $D\in\mathcal{D^{T}}$ gives me a new function $\bigcup_{t=1}^{T}\mathcal{X^{1:t-1}}\times\mathcal{Z}\rightarrow\mathcal{X}$. Analogously, fixing $\hat{f}$ at some $D\in\mathcal{D^{T}}$ also give a new function in the same function space. Since $\mathcal{D^{T}}$ is finite, we can enumerate its elements. Thus for every $D^{(i)}\in\mathcal{D^{T}}$, $i=1,\cdots,|\mathcal{D^{T}}|$, write down $\hat{f}$ fixed at $D^{(i)}$:
$\hat{f}(X^{1:t-1},D^{(i)},Z)$ $=\sum_{l=1}^{r}\phi_{l}(X^{1:t-1},Z)\beta_{l}\psi_{l1}(D^{(i)}_{1})\psi_{l2}(D^{(i)}_{2})\cdots\psi_{ld_{max}}(D^{(i)}_{d_{max}})$ (16)
Then we can express the vector field:
@@
Now consider the vector field given by the ground truth function fixed at all values of $D\in\mathcal{D^{T}}$:
$\begin{pmatrix}f(\cdot,D^{(1)},\cdot)\\ \vdots\\ f(\cdot,D^{(|\mathcal{D^{T}}|)},\cdot)\end{pmatrix}$ (18)
@@
$W^{-1}\begin{pmatrix}f(\cdot,D^{(1)},\cdot)\\ \vdots\\ f(\cdot,D^{(|\mathcal{D^{T}}|)},\cdot)\end{pmatrix}$ (19)
Since recurrent neural networks are universal approximators Schäfer and Zimmermann [2007], for every $\delta>0$ we can choose $\phi_{1},\cdots,\phi_{r}$ so that
$\text{sup}_{\mathbf{x}\in\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1},\;z\in\mathcal{Z}}\left\|\begin{pmatrix}\phi_{1}(\mathbf{x},z)\\ \vdots\\ \phi_{r}(\mathbf{x},z)\end{pmatrix}-W^{-1}\begin{pmatrix}f(\mathbf{x},D^{(1)},z)\\ \vdots\\ f(\mathbf{x},D^{(|\mathcal{D^{T}}|)},z)\end{pmatrix}\right\|^{2}_{2}\leq\delta$ (20)
@@
$\leq$ $\|W\|^{2}_{op}\delta$ (29)
So choosing $\delta=\epsilon/\|W\|^{2}_{op}$ would give us the desired result.
@@
Let $b_{n}$ be a $(T_{0_{n}}-1)\times 1$ vector, where each entry $t$ is given by $\mathbb{E}[X_{n}^{t+1}(d_{n}^{1:{t+1}})~{}|~{}x_{n}^{1:t}(d_{n}^{1:t}),z_{n}]$. Under Assumptions 1 and 2, $\beta_{n}$ is identifiable from $A_{n}$ and $b_{n}$.
@@
Let $b_{n}$ be a $(T_{0_{n}}-1)\times 1$ vector where each entry $t$ is given by $\mathbb{E}[X_{n}^{t+1}(0^{1:t+1})~{}|~{}x_{n}^{1:t}(0^{1:t}),z_{n}]$. We know from basic least-squares theory, and the given assumptions, that as long as the number $T_{0}\geq r$ of observations is larger than or equal to the dimension $r$, the system is (over-)determined and $\beta_{n}=A_{n}^{+}b$. $\Box$
@@
Assume we have a dataset of $N$ units and $T$ time points $(d_{n}^{1:T},x_{n}^{1:T},z_{n})$ generated by a model partially specified by Eq. (1). Assume also knowledge of the conditional expectations $\mathbb{E}[X_{n}^{t}(d_{n}^{1:t})~{}|~{}x_{n}^{t-1}(d_{n}^{1:t-1}),z_{n}]$ and basis functions $\phi_{l}(\cdot,\cdot)$ for all $l=1,2,\dots,r$ and all $1\leq n\leq N$ and $1\leq t\leq T$. Then, under Assumptions 1-4 applied to all individuals and all intervention levels $d$ that appear in our dataset, we have that all $\beta_{n}$ and all $\psi_{l}(d,\cdot,\cdot)$ are identifiable.
@@
Consider now the first time $t$ an intervention level $d>0$ is assigned to someone in the entire dataset. By Assumption 3, there exists a dataset $\mathcal{N}_{d}$ with the given properties where, for all $n\in\mathcal{N}_{d}$, we have $\eta_{nl}^{t}=\beta_{nl}\times\prod_{t^{\prime}=1}^{t}\psi_{l}(d_{n}^{t^{\prime}},t^{\prime},t)=\beta_{nl}$, since all interventions prior to $t$ have been at level 0.
Let $b_{dt^{\prime}}$ be a $N_{d}\times 1$ vector where each entry $i$ is given by $\mathbb{E}[X_{n_{i}}^{t+t^{\prime}}(d_{n_{i}}^{1:t+t^{\prime}})~{}|~{}x_{n_{i}}^{1:t+t^{\prime}}(d_{n_{i}}^{1:t+t^{\prime}}),z_{n_{i}}]$, for $t^{\prime}=1,2,\dots,k_{d}-1$. Consider the time-bounded case with free parameters $w_{dl,1}$, $l=1,2,\dots,r$. Given Assumption 4, we can solve for $\eta_{n}^{t+1}=A_{d_{1}}^{+}b_{d_{1}}$, the vector with entries $\eta_{nl}^{t+1}$.
@@
So far we have established identification of the $\psi(d,\cdot,\cdot)$ functions for the first intervention level $d$ that appears in the dataset (the value $d$ is not unique, as it is possible to assign different intervention levels $d^{\prime}$ in parallel at time $t$, but to different units). We now must show that, for the remaining interventions $d>0$ levels take place in our dataset after time $t$.
@@
$P_{T}(X,D)\approx Z^{-1}\sum_{n=1}^{N}w_{n}^{t^{\prime}}{\bf 1}((X,D)=(X_{n}^{t^{\prime}},D_{n}^{t^{\prime}})),\quad Z=\sum_{n=1}^{N}w_{n}^{t^{\prime}}$ (33)

### v2

@@ 6 Conclusion
We introduced an approach for predictions of sequences under hypothetically controlled actions, with a careful accounting of when extrapolation to unseen sequences of controls is warranted.
Findings. Embeddings are important given sparse categorical sequential data (Vaswani et al., 2017). However, large combinatorial interventional problems benefit from models that carefully lay down conditions for the identification of such embeddings. Naïve sequential models, however flexible, cannot fully do the heavy lifting of generalizing in the absence of structure. Information has to come from somewhere.
Limitations. Unlike the traditional synthetic control literature Abadie (2021), we assume a model for time effects based on autoregression and truncated or parametric time/intervention interactions. While it is possible to empirically evaluate the predictive abilities of the model using a validation sample, high-stakes applications (such as major interventions to counteract the effect of a pandemic) should take into consideration that uncontrolled distribution shifts may take place, and careful modeling of such shifts should be added to any analytical pipeline to avoid damaging societal implications.
Future work. Besides allowing for an explicit parameterization of drifts, accounting for unmeasured confounding that may take place among past actions and states is necessary to increase the applicability of the method. Moreover, when each action level $d$ is itself a combination of cross-sectional actions, causal energy-based models such as Bravo-Hermsdorff et al. (2023) can be combined with the compositional factorization idea to also generalize to previously unseen cross-sectional combinations of actions.
@@ Acknowledgements
We thank Zhenwen Dai and Georges Dupret from Spotify for numerous helpful comments and suggestions. JY and RS were partially funded by the EPSRC Open Fellowship The Causal Continuum: Transforming Modelling and Computation in Causal Inference, EP/W024330/1. RS acknowledges support of the UKRI AI programme, and the Engineering and Physical Sciences Research Council, for CHAI – EPSRC AI Hub for Causality in Healthcare AI with Real Data (grant number EP/Y028856/1). The authors would also like to thank the four anonymous reviewers for fruitful suggestions on how to improve the presentation of this paper.
@@ Appendix A General Background
Much of traditional causal inference aims at estimating the outcome of a single-shot treatment that is not intertwined within a sequence with uncontrolled intermediate outcomes Imbens and Rubin [2015]. However, many causal questions involve sequential treatments. For example, we might be interested in estimating the causal effects of medical treatments, lifestyle habits, employment status, marital status, occupation exposures, etc. Because these treatments may take different values for a single individual over time, they are often refereed to as time-varying treatments [Hernán and J, 2020].
In many designs, it is assumed that a time-fixed treatment applied to individuals in the population happens at the same time (i.e. there is a shared global clock for everyone). A more relaxed setting (and more practical, as described in this paper) is that (categorical) interventions can happen to different individuals at different time steps, potentially resulting in no pair of identically distributed time-series. The purpose now is to estimate the causal effect or to predict expected outcomes under time-varying treatments as a time-series, rather than estimating/imputing missing potential outcome snapshots from a fixed time interval that has taken place in the past.
One typical way of designing experiments under a binary treatment (00 or $1$) [Robins, 1986, 1997] of time-varying treatments is to have $\bar{d}_{0}=(0,0,\cdots,0)$ (“never treated”) and $\bar{d}_{1}=(1,1,\cdots,1)$ (“always treated”) as the two options in the design space. This allows us to answer what the average expected outcome contrasted between these two options. A fine-grained estimation at every time step will require further levels of contrast for other regimes, such as $d=(0,1,\cdots,1)$ and $d=(1,0,\cdots,0)$. However, the number of combinations grows exponentially on $T$, the number of time steps, even when only binary treatments are at the stake.
In this paper, we did not consider identification strategies that may be necessary when confounding is an issue. Conditional ignorability holds when we are able to block unmeasured confounders using observable confounders only. Similarly, in a time-varying setting, we achieve sequential conditional ignorability by conditioning on the time-varying covariates, including past actions (Hernán and J [2020], and Chapter 4 of Pearl [2009]). In our manuscript, we will have little to say about the uses of conditional sequential ignorability and other adjustment techniques for observational studies, as these ideas are well-understood and can be applied on top of our main results. Hence, we decided to focus on treatments that act as-if randomized so that we can explore in more detail our contributions.
@@ Appendix B Proofs
Let (i) $f(n,d^{1:T},x^{1:T},z):=f_{d}^{\mathsf{T}}(d^{1:T})f_{nxz}(n,x^{1:T},z)$, where function sequences $f_{d}(d^{1:T})$ and $f_{nxz}(n,x^{1:T},z)$ are defined for all $T\in\mathbb{N}^{+}$ and have codomain $\mathbb{R}^{r_{g}}$; (ii) $f_{dl}(d^{1:T})$ be given as in Eq. (6) with $r_{h_{l}}=1$; (iii) the $l$-th entry of $f_{nxz}(n,x^{1:T},z)$ be given by $f_{nxzl}(n,x^{1:T},z):=u_{l}(n)^{\mathsf{T}}v_{l}(x^{1:T},z)$, where both $u_{l}(\cdot)$ and $v_{l}(\cdot,\cdot)$ have codomain $\mathbb{R}^{r_{s^{\prime}}}$; (iv) $v_{l}(x^{1:t},z)$ be defined analogously to $f_{dl}(d^{1:t})$ as in Eq. (6), carrying out the fixed-size $z$ as an extra argument. Then there exists some integer $r$ and three functions $a,b$ and $c$ with codomain $\mathbb{R}^{r}$ so that $f(n,d^{1:T},x^{1:T},z)=\sum_{l=1}^{r}a_{l}(x^{1:T},z)\times b_{l}(n)\times c_{l}(d^{1:T},T)$. Moreover $c_{l}(d^{1:T},T)=\prod_{t=1}^{T}m_{l}(d^{t},t,T)$ for some function $m_{l}:\mathbb{N}^{3}\rightarrow\mathbb{R}$.
@@
$\begin{array}[]{rcl}f_{d}^{\mathsf{T}}(d^{1:T})f_{nxz}(n,x^{1:T},z)&=&\sum_{l^{\prime}=1}^{r_{s}}f_{dl^{\prime}}(d^{1:T})f_{nxzl^{\prime}}(n,x^{1:T},z)\\ \\ &=&\sum_{l^{\prime}=1}^{r_{s}}f_{dl^{\prime}}(d^{1:T})\sum_{l^{\prime\prime}}^{r_{s^{\prime}}}u_{l^{\prime}l^{\prime\prime}}(n)v_{l^{\prime}l^{\prime\prime}}(x^{1:T},z)\\ \\ &=&\sum_{l^{\prime\prime}=1}^{r_{s^{\prime}}}\sum_{l^{\prime}=1}^{r_{s}}u_{l^{\prime}l^{\prime\prime}}(n)v_{l^{\prime}l^{\prime\prime}}(x^{1:T},z)f_{dl^{\prime}}(d^{1:T})\\ \\ &=&\sum_{l=1}^{r}a_{l}(x^{1:T},z)\times b_{l}(n)\times c_{l}(d^{1:T}),\end{array}$
@@
Finally, each $c_{l}(d^{1:T})$ is by definition some $f_{dl^{\prime}}(d^{1:T})$. From Eq. (6) and the assumptions, $f_{dl^{\prime}}(d^{1:T})=g_{l^{\prime}}(d^{1:T},T,T)$, and $g_{l^{\prime}}(d^{1:t},t^{\prime},t):=h_{l^{\prime}}(g_{l}(d^{1:t^{\prime}-1},t^{\prime}-1,t),d^{t^{\prime}},t^{\prime},t)$. For $t^{\prime}>0$ and $d>0$, we have that $h_{l}^{\prime}$ is given by $h_{l^{\prime}_{1}}(g_{l^{\prime}}(d^{1:t},t^{\prime}-1,t))h_{l^{\prime}_{2}}(d^{t},t^{\prime},t)$. For $t^{\prime}=0$, $h_{l^{\prime}}(v,d,t^{\prime},t)\equiv 1$, and for $t^{\prime}>0$ and $d=0$, $h_{l^{\prime}}(v,d,t^{\prime},t)\equiv g_{l^{\prime}}(d^{1:t^{\prime}-1},t^{\prime}-1,t)$. By recursive application, this implies $f_{dl^{\prime}}(d^{1:T})=\prod_{t^{\prime}=1:d^{t^{\prime}}>0}^{T}h_{l^{\prime}_{2}}(d^{t^{\prime}},t^{\prime},T)$ and the result follows. $\Box$
@@
Given a fixed value of $T$, suppose we have a measurable function $f:\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1}\times\bigcup_{t=1}^{T}\mathcal{D}^{t}\times\mathcal{Z}\rightarrow\mathbb{R}^{d}$, $t\leq T$, where $\mathcal{X}^{1:t-1}$ and $\mathcal{Z}$ are compact, $\mathcal{D}^{t}$ is finite. Let $\mathbf{x}\in\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1}$ and let $\mathbf{d}\in\bigcup_{t=1}^{T}\mathcal{D}^{t}$. For $d=1,\cdots,d_{max}$, let $m_{d}:\bigcup_{t=1}^{T}\mathcal{D}^{t}\rightarrow\{1,\cdots,T\}$ with $m_{d}(\mathbf{d})=t\text{ if }\exists t\;s.t.\;\mathbf{d}_{t}=d\;\text{else }0$. Then, for any $\epsilon>0$, there exist $r=O(T^{d_{max}+1})$ and measurable functions $\phi_{l}:\mathcal{X}^{1:t-1}\times\mathcal{Z}\rightarrow\mathbb{R}^{d}$ and $\psi_{ld^{\prime}}:\mathcal{D}^{t}\rightarrow\mathbb{R}$, and real numbers $\beta_{l}$, such that
$\text{sup}_{\mathbf{x}\in\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1},\mathbf{d}\in\bigcup_{t=1}^{T}\mathcal{D}^{t},z\in\mathcal{Z}}\left|f(\mathbf{x},\mathbf{d},z)-\sum_{l=1}^{r}\phi_{l}(\mathbf{x},z)\times\beta_{l}\times\prod_{d=1}^{d_{max}}\psi_{ld^{\prime}}(m_{d^{\prime}}(\mathbf{d}))\right|\leq\epsilon.$ (13)
Proof. First, let
$\hat{f}(\mathbf{x},\mathbf{d},z):=\sum_{l=1}^{r}\phi_{l}(\mathbf{x},z)\times\beta_{l}\times\prod_{d=1}^{d_{max}}\psi_{ld^{\prime}}(m_{d^{\prime}}(\mathbf{d})).$
Notice that we can construct an equivalent definition of $D^{t}$ for some fixed $t$ as follows. We for now call the new object $\bar{D}^{t}$:
$\bar{\mathcal{D}}^{t}=\left\{\mathbf{v}\in\mathbb{R}^{d_{max}}\bigg{|}v_{d}\in\{0,\cdots,t\}\forall d;\;v_{d^{\prime}}\neq v_{d}\text{ whenever $d^{\prime}\neq d$ and $v_{d}\neq 0$}\right\}.$ (14)
Note that there is a one-to-one correspondence between $\bar{\mathcal{D}}^{t}$ and $\mathcal{D}^{t}$, and denote the bijection connecting them by $g:\bar{\mathcal{D}}^{t}\rightarrow\mathcal{D}^{t}$. Since both sets are finite, we can approximate $f$ with $\hat{f}$, if and only if we can approximate $f\circ i\times g\times i$ by $\hat{f}\circ i\times g\times i$. So, without loss of generality, just consider $\bar{\mathcal{D}}^{t}$ as $\mathcal{D}^{t}$.
@@
Fixing the ground truth function $f$ at $D\in\mathcal{D^{T}}$ gives us a new function $\bigcup_{t=1}^{T}\mathcal{X}^{1:t-1}\times\mathcal{Z}\rightarrow\mathcal{X}$. Analogously, fixing $\hat{f}$ at some $D\in\mathcal{D^{T}}$ also gives a new function in the same function space. Since $\mathcal{D^{T}}$ is finite, we can enumerate its elements. Thus, for every $D^{(i)}\in\mathcal{D^{T}}$, $i=1,\cdots,|\mathcal{D^{T}}|$, write down $\hat{f}$ fixed at $D^{(i)}$ as:
$\hat{f}(X^{1:t-1},D^{(i)},Z)$ $=\sum_{l=1}^{r}\phi_{l}(X^{1:t-1},Z)\beta_{l}\psi_{l1}(D^{(i)}_{1})\psi_{l2}(D^{(i)}_{2})\cdots\psi_{ld_{max}}(D^{(i)}_{d_{max}}).$ (18)
We can then express the vector field as:
@@
Now, consider the vector field given by the ground truth function fixed at all values of $D\in\mathcal{D^{T}}$:
$\begin{pmatrix}f(\cdot,D^{(1)},\cdot)\\ \vdots\\ f(\cdot,D^{(|\mathcal{D^{T}}|)},\cdot)\end{pmatrix},$ (20)
@@
$W^{-1}\begin{pmatrix}f(\cdot,D^{(1)},\cdot)\\ \vdots\\ f(\cdot,D^{(|\mathcal{D^{T}}|)},\cdot)\end{pmatrix}.$ (21)
Since recurrent neural networks are universal approximators Schäfer and Zimmermann [2007], for every $\delta>0$ we can choose $\phi_{1},\cdots,\phi_{r}$ such that
$\text{sup}_{\mathbf{x}\in\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1},\;z\in\mathcal{Z}}\left\|\begin{pmatrix}\phi_{1}(\mathbf{x},z)\\ \vdots\\ \phi_{r}(\mathbf{x},z)\end{pmatrix}-W^{-1}\begin{pmatrix}f(\mathbf{x},D^{(1)},z)\\ \vdots\\ f(\mathbf{x},D^{(|\mathcal{D^{T}}|)},z)\end{pmatrix}\right\|^{2}_{2}\leq\delta.$ (22)
@@
$\leq$ $\|W\|^{2}_{op}\delta.$ (31)
Therefore, choosing $\delta=\epsilon/\|W\|^{2}_{op}$ gives us the desired result.
@@
Let $b_{n}$ be a $(T_{0_{n}}-1)\times 1$ vector with entries $\mathbb{E}[X_{n}^{t+1}~{}|~{}x_{n}^{1:t},z_{n},\mathrm{do}(d_{n}^{1:{t+1}})]$. Under Assumptions 1 and 2, $\beta_{n}$ is identifiable from $A_{n}$ and $b_{n}$.
@@
Let $b_{n}$ be a $(T_{0_{n}}-1)\times 1$ vector where each entry $t$ is given by $\mathbb{E}[X_{n}^{t+1}~{}|~{}x_{n}^{1:t},z_{n},\mathrm{do}(0^{1:{t+1}})]$, where $0^{1:t+1}$ denotes all actions kept at the default level of 0 from time 1 to time $t+1$. We know from standard least-squares theory, and the given assumptions, that as long as the number $T_{0}\geq r$ of observations is larger than or equal to the dimension $r$ and the matrix is full rank, that the system is (over-)determined and $\beta_{n}=A_{n}^{+}b$. $\Box$
@@
Assume we have a dataset of $N$ units and $T$ time points $(d_{n}^{1:T},x_{n}^{1:T},z_{n})$ generated by a model partially specified by Eq. (1). Assume also knowledge of the conditional expectations $\mathbb{E}[X_{n}^{t}~{}|~{}x_{n}^{t-1},z_{n},\mathrm{do}(d_{n}^{1:t})]$ and basis functions $\phi_{l}(\cdot,\cdot)$ for all $l=1,2,\dots,r$ and all $1\leq n\leq N$ and $1\leq t\leq T$. Then, under Assumptions 1-4 applied to all individuals and all intervention levels $d$ that appear in our dataset, we have that all $\beta_{n}$ and all $\psi_{l}(d,\cdot,\cdot)$ are identifiable.
@@
Consider now the first time $t$ where an intervention level $d>0$ is assigned to someone in the entire dataset. By Assumption 3, there exists a dataset $\mathcal{N}_{d}$ with the given properties where, for all $n\in\mathcal{N}_{d}$, we have $\eta_{nl}^{t}=\beta_{nl}\times\prod_{t^{\prime}=1}^{t}\psi_{l}(d_{n}^{t^{\prime}},t^{\prime},t)=\beta_{nl}$, since all interventions prior to $t$ have been at level 0.
Let $b_{dt^{\prime}}$ be a $N_{d}\times 1$ vector where each entry $i$ is given by $\mathbb{E}[X_{n_{i}}^{t+t^{\prime}}~{}|~{}x_{n_{i}}^{1:t+t^{\prime}-1},z_{n_{i}},\mathrm{do}(d_{n_{i}}^{1:t+t^{\prime}})]$, for $t^{\prime}=1,2,\dots,k_{d}-1$. Consider the time-bounded case with free parameters $w_{dl,1}$, $l=1,2,\dots,r$. Given Assumption 4, we can solve for $\eta_{n}^{t+1}=A_{d_{1}}^{+}b_{d_{1}}$, the vector with entries $\eta_{nl}^{t+1}$.
@@
So far we have established identification of the $\psi(d,\cdot,\cdot)$ functions for the first intervention level $d$ that appears in the dataset (the value $d$ is not unique, as it is possible to assign different intervention levels $d^{\prime}$ in parallel at time $t$, but to different units). We now must show identification for the remaining interventions $d>0$ levels that take place in our dataset after time $t$.
@@
$P_{T}(X,D)\approx Z^{-1}\sum_{n=1}^{N}w_{n}^{t^{\prime}}{\bf 1}((X,D)=(X_{n}^{t^{\prime}},D_{n}^{t^{\prime}})),\quad Z=\sum_{n=1}^{N}w_{n}^{t^{\prime}}.$ (35)